Write the polar form of 2+i/2-i
1
Expert's answer
2020-05-18T17:41:15-0400

Let x + iy = (2 + i)/(2 - i) where 'x' represents the real component & 'y' represents the imaginary component of the given equation.


x + iy = (2 + i)/(2 - i)


Multiplying the numerator and the denominator of RHS by (2+i), we get :


x + iy = (2+i)(2+i)/(2-i)(2+i)

x + iy = (22 + i2 + 2(2)(i))/(2×2 + 2×i - (2×i) - (i×i))

We know that i2 = -1.

Therefore,

x + iy = (4 + (-1) + 4i)/(4 + 2i - 2i - (-1))

x + iy = (4 - 1 + 4i)/(4 + 1)

x + iy = (3 + 4i)/5

i.e., x + iy = (3/5) + i(4/5)


Polar form of a complex number is written as :


x + iy = r(cos"\theta" + isin"\theta")

Where, r = "\sqrt{\smash[b]{x² + y²}}" & "\theta" = arctan(y/x)


r = "\sqrt{\smash[b]{(3/5)² + (4/5)²}}" & "\theta" = arctan((4/5)/(3/5))

r = "\sqrt{\smash[b]{(9/25) + (16/25)}}" & "\theta" = arctan(4/3)

r = "\sqrt{\smash[b]{(25/25}}" & "\theta" = arctan(4/3)

r = "\sqrt{\smash[b]{1}}" & "\theta" = arctan(4/3)

"\therefore" r = 1 & "\theta" = 53.13"\degree"


Hence, the polar form of the given equation is :


(2 + i)/(2 - i) = (3/5) + i(4/5) = 1(cos(53.13"\degree") + isin(53.13"\degree"))


"\therefore" (2 + i)/(2 - i) = cos(53.13"\degree") + isin(53.13"\degree")